please help me solve this and show how you solved it so i can get a better understanding, thank you?

Solve e2x +et – 6=0.
Answer #1

Given, exted-6=0 Soth het exau so. u² tu-6-0 u² +34-24-6=0 (443) (4-2) = 0 u=-3,2 ex=-3 {Not possible? So, ex=2 x = n12) 250.in case of doubt please comment